what is capacitor

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi nitin,

A capacitor is a passive electronic component that stores energy in the form of an electrostatic field.When a voltage potential difference exists between the conductors, an electric field is present in the dielectric. This field stores energy and produces a mechanical force between the plates. In its simplest form, a capacitor consists of two conducting plates separated by an insulating material called the dielectric.

A capacitor is assumed to be self-contained and isolated, with no net electric charge and no influence from an external electric field. The conductors thus contain equal and opposite charges on their facing surfaces, and the dielectric contains an electric field.

An ideal capacitor is characterized by a single constant value, capacitance, which is measured in farads. This is the ratio of the electric charge on each conductor to the potential difference between them. In practice, the dielectric between the plates passes a small amount of leakage current. The conductors and leads introduce an equivalent series resistance and the dielectric has an electric field strength limit resulting in a breakdown voltage. capacitance C, defined as the ratio of charge ±Q on each conductor to the voltage V between them:

C = \frac{Q}{V}

The capacitance is directly proportional to the surface areas of the plates, andis inversely proportional to the separation between the plates. Capacitance also depends on the dielectric constant of the substance separating the plates.

The standard unit of capacitance is the farad, abbreviated F. Thisis a large unit; more common units are the microfarad, abbreviated µF (1 µF =10-6F) and the picofarad, abbreviated pF (1 pF = 10-12 F).